# GenAINet: Enabling Wireless Collective Intelligence via Knowledge Transfer and Reasoning

### v2

@@ III-B Multi-modal semantic reasoning
The semantic concepts are then used on agents to perform planning and reasoning, for making decisions or generating contents. CoT prompting could be used to decomposes a problem into coherent sequences serving as intermediate steps, with an interpretable window suggesting how it might reach the final solution, and gives the agent opportunity to debug when a step goes wrong [11]. CoT could be generalized to sequential or non-sequential complex structures such as tree or graph of thoughts, with rewards to optimize the path. Furthermore, the decomposed sub-tasks of a complex task can be assigned to different agents and be solved separately. Finally the outputs from each sub-task can guide agents to complete the task jointly.
Fig. 2 illustrates the process of multi-modal semantic reasoning. Semantic concepts $c$ are learned from multi-modal raw data. A semantic knowledgebase (KB) representing the semantic relations between concepts on topological structure is built. When an agent receives an input, it plans a tree of thought states $T$. In each state it retrieves concepts from semantic KB and produce an action. The final output is generated after completing the planned states. The agent can observe rewards at each state, optimize the reasoning path and the semantic concept structures in the KB.
Figure 2: Proposed pipeline of multi-modal semantic extraction, retrieval and reasoning on GenAINet agents.
@@ III-C Semantic-native GenAINet
With semantic knowledge and semantic reasoning, GenAI agents can collectively learn communication protocols and decision strategies for solving different tasks, leading to an efficient GenAINet. GenAI agents can be used for various tasks on connected devices, from generating content to making decisions. An effective communication protocol should be part of the reasoning process, guiding agents to achieve a goal, such as the accuracy of responses, the accumulated reward of actions, and so on. Furthermore, an agent should effectively manage the knowledge to transmit, memorize, and update its model according to the information freshness and energy cost.
@@
GenAINet can be implemented in a teacher-student or a distributed paradigm. The teacher-student paradigm is applicable for communication between cloud, edge, and device agents, where different sizes of LLMs are deployed. Specifically, a teacher agent with LLM trained on universal knowledge guides a student agent to perform specific tasks via knowledge transfer. To reduce communication cost, the teacher agent extracts knowledge in specific domains and transfer to the student, which is efficient for devices performing specific tasks (e.g., routing, traffic control). The distributed paradigm can be used for communication between mobile devices, machines, vehicles, where agents have similar capabilities and need to collaboratively complete tasks. Each agent has a memory of emergent knowledge which is learned from its past experience. During planning, the agent can retrieve or communicate knowledge with others to optimize decisions and update local knowledge. In a long term, the agents will share a common KB and make decisions locally, where communication costs can be minimized until observing new scenarios or tasks. Compared to standard communication systems, GenAINet is resource efficient and task effective. For example, in a remote query scenario the teacher-student paradigm can reduce load and latency in sending responses; in an autonomous vehicle (AV) scenario the distributed paradigm can improve driving safety and communication reliability.
@@
Fig. 3 illustrates an example of GenAINet communication between AVs. The car agent utilizes LLMs as knowledge retriever (sender) and decision maker (actor). Communication is part of knowledge retrieval or sharing with a remote agent, which is planned according to actions and observations in a reasoning path. This makes the information exchange efficient and effective for decision making. The retriever aggregates information from local and remote KB, pass it to the actor to generate driving commands. The actor optimizes reasoning path from environment feedback. For example, it can reward the CoT to find the best path and action. Once complete, the agent extracts knowledge (e.g., learned rules) from the recent planning and store it in the memory.
We shown here three typical use cases of semantic communication between car agents. The first case is semantic compression from multi-modal sensory data. LLMs can produce semantic latent on source textual data for lossless generation [12]. With multi-modal LLMs, semantic compression can be used to gather information efficiently from remote sensor with different field-of-views or resolutions, to improve downstream tasks such as accident detection, navigation. The second case is collaobrative knowledge, where agents exchange the abstracts and concepts from memory. This can be achieved by remote retrieval, where the agent performs semantic similarity search on the request from remote agent, and send the related information. The third case is semantic reasoning, where agents find the consequence of past actions and exchange with others. This can reduce agents’ reasoning latency and improve decision reliability. Since long-term knowledge is more generalized than short-term experience, the communication cost can be reduced.
Figure 3: Three use case of semantic native GenAINet: compression, control, and reasoning.
@@ IV Case Studies of GenAINet
In this section, we show how the GenAINet can be applied to two examples namely, wireless device query and wireless power control. Specifically, we want to show: 1) How GenAI agents can efficiently transfer knowledge to perform on-device query; 2) How GenAI agents can collaborate with reasoning to solve a wireless network problem.
IV-A Knowledge transfer for wireless device query
@@
Question answering (QnA) is a typical application of LLMs. Deploying LLMs close to the end user can largely reduce the latency and traffic burden from massive connections. Despite that light-weight LLMs are built with efficient inference techniques, they exhibit poor performance in specialized domains compared to regular LLMs. RAGs can enhance LLM’s knowledge with external database while introducing storage cost and higher latency. Envisioned from GenAINet, we propose a semantic knowledge driven on-device query. As shown in Fig. 4, In the cloud agent, its LLM extracts the related context for the queries from cloud data and compress them into latent representations to build a semantic KB. It is then sent to the LLM on a device to answer questions using RAG. The device can feedback questions to improve semantic extraction.
@@
We have performed experiments on the TeleQnA [13], including 10k Telecom domain QnA from research and standard materials. We extract the questions’ sources and build a large vector DB in the cloud. An initial batch of sampled questions are fed into GPT-3.5 to perform retrieval from the DB, under instruction prompts to generate contexts related to the questions, and build the semantic KB for RAG on device. Table I shows that the semantic KB significantly improves the TeleQnA accuracy in all categories compared to the base Llama-7B, by reducing 27% exchanged information compared to cloud based QnA. The use case shows that an LLM can extract semantic knowledge and assist other LLMs to complete a task which is more efficient than sending raw data.
Figure 4: Cloud LLMs extract semantic knowledge from raw text and send to LLM on devices for QnA with RAG.
QnA category GPT-3.5-turbo accuracy Semantic RAG accuracy Reduced bits exchange
Lexicon 96 100 31.09
Research overview 66.35 72.12 26.61
Research publication 66.98 83.72 27.43
Standard overview 64.52 85.48 26.59
Standard specification 56.38 71.28 23.91
Overall 66 80 27.13
TABLE I: QnA accuracy and reduced bits exchange (%), compared to exchanging raw QnA data.
@@ IV-B Collaborative reasoning for wireless power control
Conventional power control paradigm resorts to an optimization problem by modeling the environment and finding the best power level leading to assigned performance metric (rate, energy). With the proposed paradigm, the solving is performed by exploiting LLMs’ knowledge of wireless network (acquired during pre-training), to reason over the path from observed rate to best power. This means in particular that with GenAI agents we may release online training.
We consider a scenario of 20 paired Tx-Rx users randomly placed in a $100m^{2}$ area. Users share a common spectrum where interference is encountered. Each user pair is associated with an LLM agent, including an LLM with plan-and-solve prompting, plus a memory of past data rate observations and power allocation actions. We aim to reduce the data rate gaps towards a target so as to minimize the total power in the network. The radio environment information is unknown to the agents and we instruct LLMs to exploit its memory and its telecom knowledge. In each iteration each LLM should decide its power and explain its decision.
The mean absolute error of data rate of all users towards the target is shown in Fig. 5. We evaluated 3 scenarios: 1) standalone agents without interactions; 2) agents share memory of past observations and actions; 3) agents share an explanation of the past decisions. It can be seen that the standalone agents find the power optimization path towards the targeted rate after several rounds. With memory sharing, agents learn slowly to understand the propagation and interference environment and adjust their power in the right direction. Finally, the explanation from other agents in previous rounds can help LLMs to find better power allocation scheme quickly. The example shows that sharing reasoning outputs can effectively help LLMs to improve decision making without online training.
Figure 5: Performances of data rate gap towards target with example prompts to LLM agents on power control.

### v3

@@ III-B Multi-modal Information Fusion and Reasoning
With knowledge and information represented in corresponding semantic space, an agent can perform planning and reasoning to make decisions until the ultimate goal or target is achieved. CoT prompting could be used to decompose a unseen tasks into coherent sequences serving as intermediate (known/solvable) steps, with an interpretable window suggesting how it might reach the final solution, and gives the agent opportunity to debug when a step goes wrong [14]. CoT could be generalized to sequential or non-sequential complex structures such as tree or graph of thoughts [30, 31], with rewards to improve the reasoning path. Furthermore, the decomposed sub-tasks of a complex task can be assigned to different agents and be solved separately. Finally the outputs from each sub-task can guide agents to complete the task jointly.
Fig. 2 describes the process of multi-modal semantic reasoning. Semantics are obtained from raw data by multi-modal LLMs in a unified form. A knowledge model, where the semantic relations between knowledge are well defined in corresponding representation formats, is maintained and refined through knowledge exchange from other agents if needed. When an agent needs to execute a task passively or proactively, it plans a tree of thought states based on the semantics of the perceived information and relevant knowledge extracted from the knowledge model jointly. In each state it retrieves relevant knowledge from the knowledge model, analyze the semantics provided by the multi-model model and then produce an reasonable action. The final output is generated after completing all planned states. The agent can observe reward at each state, improve the successive reasoning path and update knowledge in the knowledge if necessary. If one reasoning path yield good performance after the full execution, LLMs can analyze, summarize and then save it into the knowledge model as new evolving knowledge.
Figure 2: Proposed pipeline of multi-modal semantic extraction, retrieval, and reasoning on GenAINet agents. Multi-modal LLMs unify information from various data modalities as semantics, enabling effective sensor data merging. Common and evolving knowledge are semantically represented (text or embeddings) to support compression, exchange, and retrieval, avoiding redundant multi-modal raw data exchange. LLMs reason and plan based on information, knowledge, and prior experiences, refining knowledge models over time. Our framework is highly adaptable to diverse tasks and heterogeneous data modalities.
@@ III-C Semantic-native GenAINet
GenAI agents can collectively learn communication protocols and decision strategies for solving different tasks, leading to an efficient GenAINet. GenAI agents can be used for various tasks on connected devices, from generating content to making decisions. An effective communication protocol should be part of the reasoning process, guiding agents to achieve a goal, such as the accuracy of responses, the accumulated reward of actions, and so on. Furthermore, an agent should effectively manage the knowledge to transmit, memorize, and update its model according to the information freshness and energy cost.
@@
There are two typical paradigms, namely, teacher-student or a distributed paradigm for GenAINet that can be implemented. The teacher-student paradigm is applicable for communication between cloud, edge, and device agents, where there is a dominant agent with considerably stronger capabilities compared to others. Specifically, a teacher agent with LLM trained on universal knowledge guides a student agent to perform specific tasks via knowledge transfer. To reduce communication cost, the teacher agent extracts and compresses knowledge with semantic equivalence in specific domains and transfer to the student, which is efficient for devices performing specific tasks (e.g., routing, traffic control). In the centralized setting, smart grids represent an excellent use case for the teacher-student paradigm in GenAINet. In such a scenario, a teacher agent, deployed at a cloud or regional control center, is responsible for managing a large-scale electricity distribution network. This agent, powered by a highly capable LLM, continuously processes data from multiple edge nodes (e.g., smart meters, substations, renewable energy sources) to perform real-time load balancing, fault detection, and predictive maintenance. To enhance efficiency, the teacher agent extracts and compresses relevant semantic knowledge, such as energy consumption patterns, fault signatures, or renewable energy forecasts. This compressed knowledge is then transmitted to student agents located at edge nodes. The student agents leverage this tailored knowledge to execute localized tasks, such as regulating energy distribution in microgrids or predicting and responding to demand spikes in specific neighborhoods. The distributed paradigm can be used for communication between mobile devices, machines, vehicles, where agents have similar capabilities and need to collaboratively complete tasks. Each agent has a memory of emergent evolving knowledge which is learned from its past experience. During planning, the agent can retrieve or communicate knowledge with others to optimize decisions and update local knowledge. In a long term, the agents will share a common knowledge model and make decisions locally, where communication costs can be minimized until observing new scenarios or tasks. Compared to standard communication systems, GenAINet is resource efficient and task effective. For example, in a remote query scenario the teacher-student paradigm can reduce load and latency in sending responses; in an Vehicle to Everything (V2X) scenario the distributed paradigm can improve driving safety and communication reliability.
@@
Fig. 3 illustrates an example of GenAINet communication between Autonomous Vehicles (AVs). The car agent utilizes LLMs as knowledge retriever (sender) and decision maker (actor). Communication is part of knowledge retrieval or sharing with a remote agent, which is planned according to actions and observations in a reasoning path. This makes the information exchange efficient and effective for decision making. The retriever aggregates information from local and remote knowledge model, pass it to the actor to generate driving commands. The actor optimizes reasoning path from environment feedback. For example, it can reward the CoT to find the best path and action. Once complete, the agent extracts knowledge (e.g., learned rules) from the recent planning and store it in the memory.
Finally, we distinguish 3 different collaboration levels that can be achieved by GenAI agents by combining the advantage of semantic communication and LLMs reasoning. The first level is the information level from multi-modal sensory data. LLMs can produce semantic latent on source textual data for lossless generation [32]. With multi-modal LLMs, semantic compression can be used to gather information efficiently from remote sensor with different field-of-views or resolutions, to improve downstream tasks such as accident detection, navigation. The second level is the decision level, where agents exchange and make their decisions jointly. This can be achieved by multi-agent debate framework, e.g., in [24, 25]. The third level is the intent/goal level, where the intent/goal of each agent is shared with other agents so that the common objective of all involved agents can be achieved jointly. For example, if the knowledge model of two agents are well synchronized, one agent merely needs to inform the another agent its intent/goal. Then the another agent can predict the decision better (even the whole plan of the former to a certain extent), leading to a huge reduction of communication costs.
Figure 3: Three levels of collaboration in GenAINet: information, decision, and goal/intent levels. At the information level, only relevant information messages are exchanged. At the decision level, agents share decisions while considering the decisions of others. At the goal/intent level, reasoning paths are generated by considering the goals of other agents, and only the goals are shared, reducing transmission overhead. For instance, car avoidance highlights how exchanging intent information rather than raw data like images suffices. These collaboration levels are adaptable to task complexity, reflecting human-like interactions in GenAINet.
@@ IV Case Studies of GenAINet
In this section, we show how the GenAINet can be applied to two examples namely, wireless device query and wireless power control. Specifically, we want to show: 1) How GenAI agents can compress semantically common knowledge in telecom context and transfer it to a wireless device to assist telecom domain knowledge query; 2) How GenAI agents can collaborate by sharing reasoning histories to solve a wireless network problem.
IV-A Semantic knowledge transfer for Telecom domain query
@@
Question answering (QnA) is a typical application of LLMs. Deploying LLMs close to the end user can largely reduce the latency and traffic burden from massive connections. Despite that light-weight LLMs are built with efficient inference techniques, they exhibit poor performance in specialized domains compared to regular LLMs. Retrieval Augmented Generations can enhance LLM’s knowledge with external database while introducing storage cost and higher latency. Envisioned from GenAINet, we propose a semantic knowledge driven on-device query. As shown in Fig. 4, in the cloud agent, its LLM extracts the related context for the queries from cloud data and compress them into embedding vectors to build a semantic Knowledge Base (KB). It is then sent to the LLM on a device to answer questions using RAG. The device can feedback questions to improve semantic extraction.
@@
We have performed experiments on the TeleQnA [33], including 10k Telecom domain QnA from research and standard materials. We extract the questions’ sources and build a large vector database in the cloud. An initial batch of sampled questions are fed into LLM to perform retrieval from the vector database, under instruction prompts to generate contexts related to the questions, and build the semantic KB for RAGs on device.
Table I shows that the semantic KB significantly improves the TeleQnA accuracy in all categories compared to both full remote knowledge and local LLMs without external knowledge. In our experiments, the proposed semantic compression scheme achieves an average compression ratio of $45\%$, significantly reducing the amount of knowledge transferred from the cloud to local devices. Despite this compression, the accuracy drop is minimal—only $5\%$—compared to querying with full knowledge, while achieving a $14\%$ improvement in accuracy compared to local LLMs without external knowledge. These results highlight the effectiveness of the semantic compression approach in balancing communication efficiency and query accuracy. Notably, the trade-off between compression ratio and accuracy varies across QnA categories, reflecting differing sensitivities of knowledge domains to compression. For instance, lexicons maintain a high accuracy of $96\%$ even with nearly $50\%$ compression, whereas research publications experience a more significant accuracy drop under similar compression levels. Queries related to standards are comparatively less sensitive to semantic compression, with only a $3\%$ accuracy drop, while lexicons and research publications exhibit larger drops of $8\%$. These observations underscore the adaptability and versatility of the proposed framework, enabling significant communication savings while maintaining competitive performance across diverse query types. This variability suggests that compression ratios can be dynamically adjusted based on the query category, paving the way for an adaptive semantic compression scheme optimized with LLMs.
QnA Category Full Knowledge (%) Compressed Knowledge (%) No Knowledge (%) Compression Ratio (%)
Lexicon 92.00 96.00 88.00 48.03
Research overview 88.46 83.65 72.12 44.84
Research publication 91.16 83.25 71.63 40.05
Standard overview 96.77 91.94 77.42 47.65
Standard specification 81.91 77.66 53.19 52.35
Overall 89.60 84.00 69.20 45.52
TABLE I: Accuracy and compression ratio (%) of LLM agents for remote query with semantic compression, compared to query with full knowledge (upper bound) and query with LLM only (lower bound).
Figure 4: Common knowledge transfer from a cloud LLM to an on-device LLM with semantic compression and Retrieval Augmented Generations (RAGs). When an on-device LLM receives a knowledge query beyond its capabilities, it first searches for its semantic vector database to find the corresponding knowledge item with maximum semantic similarity and leverage RAGs to generate accurate response. Notice that the on-device knowledge base is obtained by semantic compression of the huge cloud knowledge base.
@@ IV-B Collaborative reasoning for wireless power control
Whereas the previous case study concerns telecom language capabilities of LLM agents, the present case study concerns their ability to tune a quantity in the radio domain namely, the transmit power of a wireless device. We consider power control here because it is a very famous problem, still of practical interest whenever formulated for new settings, makes easy the task of illustrating the concepts under consideration, and is not a priori a trivial task for an LLM agent since optimizing real numbers is involved. The conventional power control paradigm consists in modeling the environment, and then finding (through optimization) the best power level leading to an assigned and fixed performance metric (data rate, energy, QoS, confidentiality, etc). With the proposed paradigm, the modeling and solving tasks are performed by the LLMs agents themselves (as performed in [34] for EV charging policies). This is performed by exploiting the vast knowledge of the LLMs about wireless networks (acquired during pre-training), to reason over the path from observed rate to best power and cooperate with other agents through semantics, e.g., in natural language. This capability in terms of flexibility is unique and has no equivalent in the vast literature of power control. This means in particular that with GenAI agents we may release online training and a wide range of performance metrics can be accommodated, depending on the wireless scenario or current network goals.
We consider a scenario of $N$ transmitter-receiver (Tx-Rx) pairs randomly placed in a $100m^{2}$ squared area. Transmitters share a common spectrum where interference occurs. The simulation assumes the absence of a central base station, reflecting scenarios in IoT settings where direct device-to-device (D2D) communication is preferred to reduce infrastructure costs and latency. Such decentralized architectures are particularly suitable for applications like smart grids or V2X networks, where devices operate autonomously and collaboratively without relying on a central controller. For simplicity, the channel gains $h_{ij}$ between every transmitter-receiver pair $(i,j)$ are taken to be realizations of Rayleigh law with a mean of $\mathbb{E}(|h_{ij}|)=1.0$. Each transmitter is initially assigned a transmit power $p_{i}$ randomly sampled from a uniform distribution over $[1,5]$ (in Watt). The system imposes a maximum power limit of $P_{\max}=10$ W. The communications are assumed to interfere each other, which translates into the signal-to-interference-plus-noise ratio (SINR) expression for Transmitter $i$: $\mathrm{SINR}_{i}=\frac{|h_{ij}|^{2}p_{i}}{1+\sum_{j\neq i}|h_{ij}|^{2}p_{j}}$. To reflect realistic transmission requirements, the target data rate for each transmitter is set as a scaled version of its achievable rate $\mu_{i}\log_{2}(1+\mathrm{SINR}_{i})$, with $\mu_{i}$ being a random scaling factor uniformly distributed over $[0.5,1.5]$. Transmitters have up to 10 rounds/iterations to adjust their power levels. The chosen reference Distributed Power Control (DPC) algorithm comes from [35]. Since the simulation setting is randomly generated, the feasibility conditions for DPC are not always satisfied, leading to complex and challenging scenarios for analysis. In the subsequent discussion, we focus on scenarios where the DPC algorithm fails to converge, highlighting the limitations and potential improvements for distributed power control using LLM-based solutions. Furthermore, each Tx-Rx pair is associated with an LLM agent, including an LLM with plan-and-solve prompting, along with a memory of its past data rate observations and power allocation actions. Our goal is to reach a given data rate target while minimizing the sum-transmit power in the network. The radio environment information is unknown to the agents, and we instruct the LLMs to exploit their memory and telecom knowledge. In each iteration, each LLM decides its power level and explains its decision. In our GenAINet framework, agents are empowered to propose, accept, or reject cooperation proposals expressed in natural language. The prompts used for individual GenAI agents and those incorporating cooperative mechanisms are shown in Fig. 5 and Fig. 6, respectively.
By defining the rate gap as the target rate minus the rate achieved after 10 rounds, Fig. 7 and Fig. 8 represent the rate gap and total power of the system for DPC, GenAI agents alone (with memory and reasoning capabilities), and the proposed GenAINet framework for $4$ Tx-Rx pairs. It can be observed that GenAI agents alone achieve a comparable rate gap to DPC without requiring online training, while maintaining significantly lower total power. Furthermore, when GenAI agents are allowed to cooperate through proposing and accepting natural language-based cooperation proposals, the GenAINet framework achieves a faster reduction in rate gap with even a slight decrease in power compared to GenAI agents alone. The increase in rate gap and total power for DPC arises from the selfish power adjustments of agents, which leads to a continual increase in interference. In contrast, GenAI agents with memory and reasoning tend to adjust their power levels more moderately to prevent drastic interference escalation. When cooperation is introduced, a subset of agents voluntarily maintain or slightly decrease their power levels, contributing to improved overall system performance.
We further compare the performance of different approaches for varying numbers of Tx-Rx pairs in Table II. The proposed GenAINet approach consistently outperforms other approaches, except in the power level for the $2$-transmitter case. This observation suggests that cooperation may not be necessary when the number of transmitters is small. Additionally, the amount of exchanged information, measured as the number of messages sent per transmitter, is recorded in Table II. On average, transmitters sent approximately $4$ to $5$ messages during the $10$-turn game, regardless of the number of Tx-Rx pairs, highlighting the scalability of the framework. It is worth noting that no prior knowledge of power control, online learning, or task-specific fine-tuning was required for the LLMs in our simulations. This example demonstrates the zero-shot adaptation and generalization capabilities of LLMs, effectively compensating for the potential high resource consumption of LLM-based solutions for complicated reasoning tasks.
{mdframed}
[linecolor=black, linewidth=2pt, roundcorner=10pt] System prompt: Consider a wireless network with {user_num} transmitter-receiver pair sharing the same spectrum. Each user has allocated bandwidth of {bandwidth} kHz. Maximum power allowed is Pmax. The target rate of all Tx-Rx pairs: rate kbps.
User prompt: You are {user_id}. Your current transmission power: {power} W. Your current transmission rate: {rate} kbps. You should adjust your transmit power to reach the targeted rate: {target} kbps while help to minimize the total power of the network. Think step by step, consider interference with others and your past history. Once you made a final decision, output in the following format: {action:power, explanation:thought}.
Memory: {user:user_id, round:round_num, observation:rate, action:power, explanation:thought}
Figure 5: Prompt template for collaborative power control leveraging GenAINet with memory and reasoning. The target of all agents is shared in the beginning then only each agent has accessed to its local memory and leverage it reasoning capabilities.
{mdframed}
[linecolor=black, linewidth=2pt, roundcorner=10pt] System prompt: Consider a wireless network with {user_num} paired Tx and Rx users sharing the same spectrum. Each user has allocated bandwidth of {bandwidth} kHz. Maximum power allowed is Pmax. The target rate of all users: rate kbps.
User prompt: You are {user_id}. Your current transmission power: {power} W. Your current transmission rate: {rate} kbps. You should adjust your transmit power to reach the targeted rate: {target} kbps while help to minimize the total power of the network. Think step by step, consider interference and cooperation proposals with others. Be careful of not increasing drastically your power since it is damaging to increase the interference level for all Tx-Rx pairs. If you received proposals, your explanation should include the corresponding reasoning for accepting or rejecting them. You can propose cooperation plan to other users. If cooperation is needed, send a concise proposal to another user using this format: “To User [id]: [content of proposal]”. Once you made a final decision, output in the following format: {action:power, message:proposal explanation:thought}.
Memory: {user:user_id, round:round_num, observation:rate, action:power, message:proposal, explanation:thought}
Figure 6: Prompt template for collaborative power control leveraging GenAINet with communication. The target of all agents is shared in the beginning, then each agent has accessed to its local memory, leverages it reasoning capabilities and sent/receive cooperation proposal. Agents are allowed to accept or rejected received proposals, leading to the a flexible and adaptive framework.
Figure 7: Absolute rate gap for the 4-user case versus the number of rounds for the distributed power control algorithm [34], the standalone approach where GenAI agents utilize memory and LLM reasoning, and the proposed GenAINet framework with natural language-based cooperation. Cooperation between GenAI agents results in faster convergence.
Figure 8: Total power for the 4-user case versus the number of rounds is presented for the distributed power control algorithm [34], the standalone approach where agents utilize memory and LLM reasoning, and the proposed GenAINet framework with natural language-based cooperation. Cooperation between GenAI agents leads to much smaller total power for preventing interference from growing rapidly.